# NeuralOM: Neural Ocean Model for Subseasonal-to-Seasonal Simulation

### v2

@@ Abstract
Accurate Subseasonal-to-Seasonal (S2S) ocean simulation is critically important for marine research, yet remains challenging due to its substantial thermal inertia and extended time delay. Machine learning (ML)-based models have demonstrated significant advancements in simulation accuracy and computational efficiency compared to traditional numerical methods. Nevertheless, a significant limitation of current ML models for S2S ocean simulation is their inadequate incorporation of physical consistency and the slow-changing properties of the ocean system. In this work, we propose a neural ocean model (NeuralOM) for S2S ocean simulation with a multi-scale interactive graph neural network to emulate diverse physical phenomena associated with ocean systems effectively. Specifically, we propose a multi-stage framework tailored to model the ocean’s slowly changing nature. Additionally, we introduce a multi-scale interactive messaging module to capture complex dynamical behaviors, such as gradient changes and multiplicative coupling relationships inherent in ocean dynamics. Extensive experimental evaluations confirm that our proposed NeuralOM outperforms state-of-the-art models in S2S and extreme event simulation. The codes are available at https://github.com/YuanGao-YG/NeuralOM.
1 Introduction
Subseasonal-to-Seasonal (S2S) simulation represents a simulation of a dynamic system 2 to 6 weeks in advance liu2025cirt . Accurate S2S ocean simulation plays a pivotal role in marine research and activities, yet it remains a significant challenge due to the ocean’s substantial thermal inertia and prolonged time delay subramanian2019ocean ; lee2025observed . Traditional numerical ocean modeling systems, such as the Mercator Ocean Physical System (PSY4) and Real-Time Ocean Forecast System (RTOFS), rely on physics-driven numerical models to simulate ocean phenomena. These models use supercomputers to solve complex partial differential equations based on current ocean conditions francis2020high . Although these numerical systems achieve high accuracy, they are computationally expensive and time-consuming, requiring hours of computation even with hundreds of computational nodes. Moreover, enhancing the accuracy of these numerical methods is highly challenging, as they depend significantly on human expertise in understanding the complex physical laws governing the ocean environment.
Machine Learning (ML) provides a new perspective for ocean simulation due to several orders of magnitude faster than traditional numerical methods and high efficiency liu2024foundation  hao2025deep . Some spatio-temporal prediction algorithms, such as NMO wu2024neural , neXtSIM finn2024generative , OceanVP shi2024oceanvp focus on regional ocean modeling. And some ML-based task-specific ocean models are proposed, such as ENSO (El Niño/Southern Oscillation) prediction ham2019deep  chen2025combined , MJO (Madden Julian Oscillation) prediction kim2021deep  delaunay2022interpretable  shin2024deep , and MHWs (Marine Heatwaves) prediction shi2024physics  jacox2022global  shu2025advanced . In recent years, learning from the algorithms of ML-based global atmosphere foundation models pathak2022FourCastNet  bi2023accurate , ML-based global ocean foundation models have emerged, such as AI-GOMS xiong2023ai , Xihe wang2024xihe , and WenHai cui2025forecasting . Some foundation models achieve better performance than traditional numerical models. However, ML-based global ocean foundation models still face several core challenges: ❶ Slow-changing processes are hard to model. Due to substantial thermal inertia and extended time delay of the ocean system, it is difficult to learn the small changes effectively. ❷ Subseasonal-to-seasonal simulation lacks of adaptive modeling. Different time scales incorporate multi-scale ocean movements. Focusing on single-scale movement is hard to model consistently with the evolution process of ocean dynamics.
To address these challenges, this work proposes NeuralOM, a neural ocean model for subseasonal-to-seasonal simulation. Specially, to model the ocean’s slowly changing nature, we propose an ocean-specific multi-stage framework to learn large-scale to small-scale characteristics step by step. Under this framework, variables with seasonal periodicity, such as sea salinity, sea temperature, and sea surface height, are first subtracted from the climatology so that the model can better capture subtle changes. Additionally, we introduce a multi-scale interactive messaging module to capture complex dynamical behaviors, such as gradient changes and multiplicative coupling relationships inherent in ocean dynamics, which contributes to the S2S time scale simulation by adaptively modeling multi-scale movements.
The contribution of this work can be summarized as: (1) Stable Subseasonal-to-Seasonal Ocean Simulation Framework. We proposed an ocean-specific multi-stage framework tailored for the slow change nature of the ocean dynamic system, which achieves stale ocean simulation at S2S time scale. (2) Adaptively Modeling of Multi-scale Movements for Ocean Dynamic System. We propose a multi-scale interactive graph neural network to adaptively model the evolution process of ocean dynamic system. (3) Competitive Experimental Results. Extensive experiments demonstrate that our proposed model outperforms state-of-the-art models in S2S and extreme event simulation.
2 Related Work
ML-based Task-specific Ocean Model. The goal for this type of ocean model is to simulate specific ocean phenomena. Recent achievements includes ENSO (El Niño/Southern Oscillation) pediction ham2019deep  hu2021deep  chang2023predicting  lyu2024resonet  chen2025combined , MJO (Madden Julian Oscillation) prediction kim2021deep  delaunay2022interpretable  yang2023exploring  shin2024data  shin2024deep , MHWs (Marine Heatwaves) prediction jacox2022global  sun2024deep  parasyris2025marine  shu2025advanced , Sea Surface Height anomaly (SSHa) prediction song2020deep  liu2020sea . These models have demonstrated promising performance in comparison to traditional methods. However, several models overlook the interactions between multiple variables, suggesting that ML-based ocean models could potentially achieve enhanced accuracy by incorporating multi-variable interactions across 3D depth levels with a foundation model.
ML-based Global Ocean Foundation Model. The global ocean foundation models have significant achievements with higher accuracy and better efficiency than traditional numerical models. AI-GOMS xiong2023ai is an ocean foundation model that employs MAE he2022masked to simulate a multi-layer ocean state. Xihe wang2024xihe achieves an eddy-resolving simulation based on the hierarchical transformer. Some research, such as ORCA-DL guo2024data , OCE subel2024building , and DLESyM cresswell2024deep , targets decadal to climate-scale simulation, which indicates tremendous potential of ML-based models in climate simulation. LangYa yang2024langya is a cross-spatiotemporal ocean model with large-language model-based time embedding. Subsequently, WenHai cui2025forecasting incorporates domain knowledge into Swin Transformer liu2021swin and achieves better performance than the numerical model.
3 Method
Problem Setup In this work, we formulate ocean simulation as an autoregressive problem. For each time step $t$, we use the ocean state and forcing to simulate the next time step ocean state. The ocean state is composed of ocean depth level variables $\mathbf{O}^{d}_{t}$ and surface variables $\mathbf{O}^{s}_{t}$, while the forcing includes atmosphere surface variables $\mathbf{F}_{t}$ and $\mathbf{F}_{t+1}$. These components are concatenated along the channel dimension to form the combined input: $\mathbf{X}_{t}=[\mathbf{O}^{d}_{t},\mathbf{O}^{s}_{t},\mathbf{F}_{t},\mathbf{F}_{t+1}]\in\mathbb{R}^{N\times d}$, where $N=H\times W$ represents the number of grid locations, and $d=d_{d}+d_{s}+d_{f}$ is the total number of input variables. Here, $d_{d}$, $d_{s}$, and $d_{f}$ are the numbers of ocean depth level variables, surface level variables, and forcing variables, respectively. In our setup, the input contains 101 variables: 92 depth level variables, 1 surface level variable, and 8 forcing variables. The model aims to simulate the next time ocean state $\hat{\mathbf{O}}_{t+1}$, given the current input $\mathbf{X}_{t}$, capturing the spatiotemporal evolution of the ocean: $\hat{\mathbf{O}}_{t+1}=\text{Model}(\mathbf{X}_{t};\Theta),$ where $\Theta$ denotes the model parameters. $\hat{\mathbf{O}}_{t+1}=[\hat{\mathbf{O}}^{d}_{t+1},\hat{\mathbf{O}}^{s}_{t+1}]\in\mathbb{R}^{N\times d_{o}}$, where $d_{o}=d_{d}+d_{s}$ is the total number of ocean variables. The training objective is to minimize the relative mean squared error (MSE) between the simulations and the true values across all time steps: $\min_{\Theta}\frac{1}{T}\sum_{t=0}^{T-1}\frac{\left\|\hat{\mathbf{O}}_{t+1}-\mathbf{O}_{t+1}\right\|_{2}^{2}}{\left\|\mathbf{O}_{t+1}\right\|_{2}^{2}}$. During inference, we employ a rollout strategy to simulate longer sequences. Starting from the initial state $\mathbf{X}_{0}$, the model recursively uses its previous simulations as the next input: $\hat{\mathbf{O}}_{t+1}=\text{Model}(\hat{\mathbf{X}}_{t};\Theta),\quad t=0,1,2,\dots,T-1.$
3.1 Ocean-specific Multi-stage Framework
Due to its substantial thermal inertia and extended time delay, the ocean changes slowly compared to its original state, making it difficult to effectively learn and capture changes. Therefore, as depicted in Figure 1(a), we propose an ocean-specific multi-stage framework (OMF) to learn small change nature. Specifically, different from previous works such as AI-GOMS, Xihe, and WenHai, we first remove the climatological mean from variables exhibiting strong periodicity before feeding them into the network, thereby directing the model’s attention toward small anomalies:
$O_{t}^{periodic}={O_{t}}^{original}-{O_{t}}^{climatology},$ (1)
where, $O_{t}^{periodic}$ denotes periodic ocean variables (sea salinity, sea temperature, and sea surface height), ${O_{t}}^{original}$ represents the original ocean variables, ${O_{t}}^{climatology}$ is the climatology calculated by the historical data from 1993 to 2017, exclude from valid and test data. Therefore, the input ocean variables ${O}_{t}=[{O}^{periodic}_{t},{O}^{u,v}_{t}]\in\mathbb{R}^{N\times d_{o}}$. ${O}^{u,v}_{t}$ is sea stream velocity. Then, inspired by mojgani2024interpretable that intends to reduce step error for scientific data prediction and a multi-step supervised method trask2022hierarchical , we adopt a stage-wise learning strategy to sequentially capture the ocean’s gradual evolution. It is worth noting that although the previous works could reduce single-step error for some ideal datasets, they are hard to achieve long-term simulation in complex real-world ocean scenes. In contrast, our proposed OMF focuses on long-term simulation and achieves stable subseasonal-to-seasonal ocean simulation. Specifically, the intermediate results of the next step ocean state $\hat{O}_{t+1}$ can be acquired using the combined input $X_{t}=[O_{t},F_{t},F_{t+1}]\in\mathbb{R}^{N\times d}$:
$\hat{O}^{1}_{t+1}={\mathcal{F}^{M}_{1}}(X_{t}),$ (2)
where, ${\mathcal{F}^{M}_{1}}$ is the first model with M-step finetune, which can be trained step by step from the 1-step supervised model. Subsequently, as shown in Figure 1(a), we apply the second model to learn the residue between $\hat{O}^{1}_{t+1}$ and ${O}_{t+1}$, thereby getting the intermediate results produced by the second model:
$\hat{O}^{2}_{t+1}={\mathcal{F}^{N}_{2}}(\hat{O}^{1}_{t+1})+\hat{O}^{1}_{t+1},$ (3)
where, ${\mathcal{F}^{M}_{2}}$ is the second model with N-step finetune, which is trained step by step. For the last model:
$\hat{O}_{t+1}={\mathcal{F}^{N}_{Q}}(\hat{O}^{Q-1}_{t+1})+\hat{O}^{Q-1}_{t+1},$ (4)
where, ${\mathcal{F}^{N}_{Q}}$ denotes the Qth model with N-step finetune. In our experiment, we set M=6, N=10, and Q=2 as an example to demonstrate that the proposed ocean-specific multi-stage framework could improve the long-term simulation for complex ocean dynamic systems.
3.2 Multi-scale Interactive Graph Neural Networks
3.2.1 Graph Encoder
Following the previous works lam2023learning  gao2025oneforecast , we employ the multi-scale graph encoder to map the data from lat-lon grids into a graph structure. The multi-scale graph structure can be defined as:
$MLP=\mathcal{N}(\sigma(\mathcal{L}(\cdot))),$ (7)
where, $\mathrm{ESMLP}(\cdot)$ is the Edge Sum MLP pfaff2020learning . The mesh node features are updated by an MLP:
@@ 3.2.1 Graph Encoder
Figure 1: Overview of Our NeuralOM. (a) The overall architecture of the ocean-specific multi-stage framework, input variables (subtracting climatology for periodic variables), multi-model stage, and visualization of simulation ocean variables; (b) The proposed multi-scale interactive graph neural networks; (c) The global ocean simulation module uses a rollout approach to generate future results.
3.2.2 Multi-scale Interactive Massaging
To make the message passing more consistent with the evolution process of the ocean dynamics system, we propose a module called Multi-scale Interactive Messaging (MIM). MIM includes a physical process-based edge update module and a multi-scale node update module. In our setup, 16 MIMs are applied to conduct massaging.
Physical Process-based Edge Update Module. In ocean simulation, ocean currents, temperature, salinity, and other factors interact with each other, and these interactions span different scales and layers. Unlike traditional MLP-based update strategy, which update edges by simply concatenating edges and nodes, we introduce a physical process-based edge update module to more accurately simulate the dynamic relationships in the ocean system. For the sender node feature $\mathbf{h}_{s(i)}$ and receiver nodes feature $\mathbf{h}_{r(i)}$, we first calculate their mutual interaction:
$\mathbf{h}_{d(i)}=\mathbf{h}_{s(i)}-\mathbf{h}_{r(i)},$ (12)
$\mathbf{h}_{mp(i)}=\mathbf{h}_{s(i)}*\mathbf{h}_{r(i)},$ (13)
$\mathbf{h}_{cos(i)}=cos(\mathbf{h}_{s(i)},\mathbf{h}_{r(i)}),$ (14)
$\mathbf{h}_{s(i)}=\mathcal{R}(\mathbf{h}_{s(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{cos(i)}),\mathbf{h}_{r(i)}=\mathcal{R}(\mathbf{h}_{r(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{mp(i)},\mathbf{h}_{cos(i)}).$ (16)
${\mathcal{E}_{f}}={\mathcal{E}_{f}}^{\prime}+{\mathcal{E}_{f}}.$ (20)
Multi-scale Node Update Module. Unlike traditional node update strategies, which typically sum or average the features of neighboring nodes, we consider the movements of the ocean at different scales. To achieve adaptive adjustment of aggregation methods across different scales, we introduce a multi-scale node update module that dynamically adjusts the weights of the aggregation strategy. For rapid local ocean dynamics with smaller scales, the gating mechanism automatically tends to favor sum aggregation, capturing the local rapid dynamic changes quickly and accurately. Conversely, when the model identifies that the motion scale around the node is larger and the changes are slower, it tends to prefer mean aggregation, which achieves a smoother and more robust representation of the large-scale ocean state.
$h_{sum}^{\prime}=\mathrm{MLP}_{sum}(h,\sum\mathcal{E}_{f}^{\prime}).$ (21)
$h_{mean}^{\prime}=\mathrm{MLP}_{mean}(h,\sum\mathcal{E}_{f}^{\prime}).$ (22)
3.2.3 Decoder
$\mathcal{V}_{f}^{G}=\mathcal{V}_{f}^{G}+\mathcal{V}_{f}^{G^{\prime}}.$ (27)
Finally, we use an MLP to predict the next step results:
$\hat{O}_{t+1}=\operatorname{MLP}_{d2}\left(\mathcal{V}_{f}^{G}\right).$ (28)
3.3 Optimization

### v3

@@ Abstract
Long-term, high-fidelity simulation of slow-changing physical systems, such as the ocean and climate, presents a fundamental challenge in scientific computing. Traditional autoregressive machine learning models often fail in these tasks as minor errors accumulate and lead to rapid forecast degradation. To address this problem, we propose NeuralOM, a general neural operator framework designed for simulating complex, slow-changing dynamics. NeuralOM’s core consists of two key innovations: (1) a Progressive Residual Correction Framework that decomposes the forecasting task into a series of fine-grained refinement steps, effectively suppressing long-term error accumulation; and (2) a Physics-Guided Graph Network whose built-in adaptive messaging mechanism explicitly models multi-scale physical interactions, such as gradient-driven flows and multiplicative couplings, thereby enhancing physical consistency while maintaining computational efficiency. We validate NeuralOM on the challenging task of global Subseasonal-to-Seasonal (S2S) ocean simulation. Extensive experiments demonstrate that NeuralOM not only surpasses state-of-the-art models in forecast accuracy and long-term stability, but also excels in simulating extreme events. For instance, at a 60-day lead time, NeuralOM achieves a 13.3% lower RMSE compared to the best-performing baseline, offering a stable, efficient, and physically-aware paradigm for data-driven scientific computing. Code link: https://github.com/YuanGao-YG/NeuralOM.
Introduction
The long-term simulation of complex physical systems (Wang et al. 2024b), such as the global ocean and climate, is a cornerstone of modern science yet remains a great challenge. While machine learning (ML) models offer a computationally efficient path to complement traditional numerical solvers, they face a fundamental bottleneck in such tasks: long-term stability. This issue is especially severe in slow-changing systems, where the state evolves in a subtle manner. In standard autoregressive frameworks, minor prediction errors at each step inevitably accumulate, causing the simulation to diverge catastrophically from the true physical dynamics, a well-known problem of compounding error. This error cascade often leads data-driven models to fail in long-term rollouts. As shown in Figure 1 (bottom left), a state-of-the-art baseline model generates substantial, physically unrealistic artifacts in its long-range forecast.
Figure 1: NeuralOM overcomes the long-term instability challenge in simulating slow-changing physical systems. Shown is a 60-day Subseasonal-to-Seasonal (S2S) simulation for sea surface temperature anomaly. (Bottom Left) A state-of-the-art baseline (CirT) collapses due to compounding errors, producing physically unrealistic artifacts and failing to capture the ocean state. (Bottom Right) In contrast, NeuralOM maintains high fidelity and long-term stability, accurately reproducing the complex patterns of the ground-truth, as highlighted in the detailed view (red dashed lines).
To address this challenge, existing machine learning approaches have been explored, but they are limited in two key aspects. ❶. First, they often struggle to capture the fine-grained dynamics inherent to slow-changing systems. Many models are designed to predict sharp, significant shifts, making them ill-suited for learning the small, incremental updates that define ocean dynamics. This typically results in either over-smoothing the predictions or failing to capture critical physical anomalies. ❷. Second, the internal mechanisms of these models are often physics-agnostic. For instance, standard message passing in Graph Neural Networks (GNNs) (Keisler 2022) treats all node interactions uniformly, unable to distinguish between fundamentally different physical processes like gradient-driven flows and multiplicative couplings between variables. Furthermore, they lack the ability to adaptively model phenomena across different spatio-temporal scales, from large-scale, slow-moving currents to small-scale, energetic eddies.
To systematically address these issues, we introduce NeuralOM, a neural operator framework designed for stable, high-fidelity simulation of slow-changing dynamics. NeuralOM’s core is built on two key innovations. First, we propose a Progressive Residual Correction Framework. This framework decomposes the complex forecasting task into a cascade of refinement steps, where each step focuses on learning and correcting the residual error from the previous one. This strategy of ”small, iterative corrections” enables the model to capture extremely fine-grained dynamic changes, thereby effectively suppressing long-term error accumulation. Second, we design a novel Physics-Guided Graph Network with a built-in adaptive messaging mechanism. It incorporates physical priors into the network by explicitly modeling gradient-driven flows and multiplicative couplings between variables. Crucially, its aggregation scheme is not fixed but dynamically adjusts, allowing it to inherently handle multi-scale dynamics.
We validate NeuralOM on the challenging benchmark of global Subseasonal-to-Seasonal (S2S) ocean simulation and forecasting. Extensive experiments provide strong evidence of its superior performance:
Experimental Observations. ❶ State-of-the-art accuracy in long-term simulation: As shown in Table 1, in $60$-day simulation, NeuralOM achieves a normalized RMSE of just $0.7015$, a 13.313.3% reduction compared to the best-performing baseline, WenHai ($0.8093$). It also improves the Anomaly Correlation Coefficient (ACC) by 12.012.0%, indicating more accurate tracking of the physical field’s evolution. ❷ Superior visual fidelity and physical consistency: As illustrated in Figure 1 and Figure 3, unlike baseline models like CirT that collapse due to error explosion, NeuralOM produces physically consistent forecasts that remain highly aligned with the ground truth even at $60$ days, successfully preserving critical fine-scale structures. ❸ Validated effectiveness of our designs through ablation studies: As detailed in Table 3 and Table 4, removing any of our core components, either the Progressive Residual Correction Framework or the Physics-Guided Adaptive Graph Messaging, leads to a significant drop in performance. In particular, incorporating climatology priors improves the ACC metric by more than $5$-fold, highlighting that every design choice is crucial to achieve accurate and stable simulations.
In summary, our main contributions are threefold: (1) a progressive residual correction framework designed for slow-changing systems; (2) a physics-guided graph network with an adaptive, multi-scale messaging mechanism; and (3) state-of-the-art performance on the S2S ocean simulation benchmark, establishing a new paradigm for data-driven scientific computing.
Related Work
In recent years, previous works in spatio-temporal mining have achieved stupendous achievements (Pfaff et al. 2020; Li et al. 2020; Wu et al. 2023, 2024d; Wang et al. 2024a; Wu et al. 2024b, a). However, many of these methods are only applicable in idealized or relative simple scenario, whereas ocean system is a complex dynamical system in the real world, requiring the design of specialized methods for accurate prediction (Wu et al. 2024c). Research in data-driven ocean simulation is primarily divided into two streams: task-specific predictive models and general-purpose global foundation models.
Task-Specific ML Models.
These models aim to simulate specific oceanic or atmospheric phenomena and have achieved impressive results in their respective domains. Examples include work on El Nio-Southern Oscillation (ENSO) prediction (Ham, Kim, and Luo 2019; Hu et al. 2021; Chang et al. 2023; Lyu et al. 2024; Chen et al. 2025), as well as models for the Madden-Julian Oscillation (MJO) (Kim et al. 2021; Delaunay and Christensen 2022; Yang et al. 2023; Shin et al. 2024a, b) and Marine Heatwaves (MHWs) (Jacox et al. 2022; Sun, Jing, and Liu 2024; Parasyris et al. 2025; Shu et al. 2025). The success of these models validates the immense potential of machine learning for complex geoscience problems. However, their limitation is also apparent: they are typically highly optimized for a single objective and often overlook the complex interactions between different physical variables and across multiple scales. This makes them difficult to generalize into a foundation model capable of capturing the full dynamics of the ocean system.
Global Ocean System Foundation Models.
Inspired by the success of foundation models in atmospheric science (Pathak et al. 2022; Bi et al. 2023; Chen et al. 2023a, b; Lam et al. 2023; Oskarsson et al. 2024; Nguyen et al. 2024; Kochkov et al. 2024; Gao et al. 2025), a series of global ocean foundation models have recently emerged, such as AI-GOMS (Xiong et al. 2023), Xihe (Wang et al. 2024c), LangYa (Yang et al. 2024), and WenHai (Cui et al. 2025). These models aim to build a unified digital twin of the ocean and have already surpassed traditional numerical methods in computational efficiency and, in some cases, accuracy (Hao et al. 2025). However, they still face the two core challenges we identified in the introduction. First, they generally adopt standard autoregressive frameworks and have not fundamentally solved the problem of error accumulation in long-term simulations, leading to a decline in forecast fidelity over time. Second, their designs often lack sufficient physical constraints. For example, while models like WenHai (Cui et al. 2025) incorporate domain knowledge into a Swin Transformer (Liu et al. 2021), their feature interaction mechanisms remain relatively physics-agnostic, failing to explicitly model physical processes like gradient changes or variable couplings, nor can they adaptively handle multi-scale dynamics.
@@ Introduction
Figure 2: Overview of Our NeuralOM. (a) The overall architecture of the progressive residual correction framework, input variables (subtracting climatology for periodic variables), progressive residual correction stage, and visualization of simulation ocean variables; (b) The proposed physics-guided graph network; (c) The global ocean simulation module uses a rollout approach to generate future results.
NeuralOM is designed precisely to address these limitations. Unlike task-specific models, it is a general-purpose global model. Compared to existing global foundation models, NeuralOM introduces two fundamental breakthroughs: (1) Its Progressive Residual Correction Framework is a novel architecture specifically designed to suppress error accumulation in slow-changing systems and ensure long-term stability. (2) Its Physics-Guided Graph Network, with its adaptive messaging mechanism, explicitly integrates physical interactions (e.g., gradients and couplings) into the model, addressing the physics-agnostic problem of existing models.
Methodology
Preliminaries.
We formulate the ocean system simulation as an autoregressive forecasting problem. Given the system state $\mathbf{O}_{t}$ at the current time step $t$ (such as ocean temperature and salinity) and the external forcing $\mathbf{F}_{t}$, $\mathbf{F}_{t+1}$ (e.g., atmospheric fields), our goal is to predict the system state at the next time step, $\hat{\mathbf{O}}_{t+1}$. This process is represented by a neural network $\mathcal{M}$:
$\hat{\mathbf{O}}_{t+1}=\mathcal{M}(\mathbf{O}_{t},\mathbf{F}_{t},\mathbf{F}_{t+1};\Theta),$ (1)
where $\Theta$ represents the learnable parameters of the model. During inference, the model performs long-range forecasting by recursively feeding its own prediction from the previous step, $\hat{\mathbf{O}}_{t}$, as input for the next step, starting from an initial state $\mathbf{O}_{0}$. Further details on the specific variables and data preprocessing used in our experiments are provided in the Appendix.
Progressive Correction for Stability
To overcome the challenges of error accumulation and the physics-agnostic nature of existing models, we propose NeuralOM, a stable and efficient simulation framework designed for slow-changing physical systems. Its core design comprises a Progressive Residual Correction Framework and a Physics-Guided Graph Network that serves as its core engine.
Progressive Residual Correction
When simulating slow-changing systems, directly predicting the entire next state $\mathbf{O}_{t+1}$ is difficult, as the true change ($\mathbf{O}_{t+1}-\mathbf{O}_{t}$) is remarkably small. Most models struggle to capture this subtle signal accurately, leading to rapid error accumulation.
To address this, we design a Progressive Residual Correction framework. It decomposes the complex single-step prediction task into a series of simpler, cascaded correction steps. First, a base model makes an initial, coarse prediction of the next state. Subsequently, a series of residual models are invoked, with each model focusing on learning and correcting the prediction error from the previous stage. This process can be abstractly formulated as:
$\hat{\mathbf{O}}_{t+1}^{(0)}$ $=\mathcal{M}_{\text{base}}(\mathbf{X}_{t})$ (2)
$\hat{\mathbf{O}}_{t+1}^{(q)}$ $=\hat{\mathbf{O}}_{t+1}^{(q-1)}+\mathcal{M}_{\text{residual}}^{(q)}(\hat{\mathbf{O}}_{t+1}^{(q-1)}),\quad\text{for }q=1,\dots,Q,$ (3)
where $\mathbf{X}_{t}$ is the input of model, $\hat{\mathbf{O}}_{t+1}^{(0)}$ is the initial prediction, and $\hat{\mathbf{O}}_{t+1}^{(q)}$ is the refined prediction after $q$ correction stages. The final output is $\hat{\mathbf{O}}_{t+1}=\hat{\mathbf{O}}_{t+1}^{(Q)}$. This approach shifts the model’s focus from predicting a large state vector to capturing fine-grained, physically meaningful anomalies, thereby significantly enhancing long-term stability and fidelity. Additionally, for variables with strong periodicity (e.g., sea salinity, sea temperature, and sea surface height), we first subtract their climatological mean before feeding them into the model to further help it focus on learning small anomaly changes.
Physics-Guided Adaptive Graph Messaging.
Each model within this framework ($\mathcal{M}_{\text{base}},\mathcal{M}_{\text{residual}}$) is realized by our novel Physics-Guided Graph Network. Unlike standard GNNs, our model deeply integrates physical priors into the graph’s message passing and node update procedures through its core Physics-guided Adaptive Graph Messaging module.
Physics-Inspired Edge-Interactions.
Traditional GNNs are often physics-agnostic, typically just concatenating adjacent node features and passing them through an MLP. To more realistically emulate the dynamic relationships between physical variables, we design physics-inspired edge interactions (PEI). The message between two nodes (i.e., the edge feature) is constructed not blindly, but through a set of explicit physical operators, including: (1) feature differencing to model gradient-driven flows; (2) multiplicative coupling to capture the joint effects of variables (like temperature and salinity on density); and (3) cosine similarity to identify similar water masses. These operators capture core relationships in ocean dynamics and provide a strong inductive bias. This process can be abstractly represented as:
$\mathbf{e}_{uv}=f_{\text{PEI}}(h_{u},h_{v})=\phi([\underbrace{h_{u}-h_{v}}_{\text{Gradient}},\underbrace{h_{u}\odot h_{v}}_{\text{Coupling}},\underbrace{\text{sim}(h_{u},h_{v})}_{\text{Similarity}}]),$ (4)
where $h_{u},h_{v}$ are node features and $\phi$ is a neural network that fuses these physical features.
Multi-scale Adaptive Node Aggregation.
Physical systems often contain motions at different scales. To capture both large-scale, slow-changing processes and small-scale, energetic events simultaneously, we introduce a dynamic gating mechanism for aggregating neighbor information. This mechanism adaptively arbitrates between sum and mean aggregation based on local dynamics. For localized, high-energy events (like eddies), the gate favors sum aggregation to preserve their intensity. For large-scale, smooth flows, it favors mean aggregation to obtain a more robust, macroscopic representation.
$h^{\prime}_{v}=(1-\gamma)\cdot\text{Aggregate}_{\text{mean}}(\{\mathbf{e}_{uv}\})+\gamma\cdot\text{Aggregate}_{\text{sum}}(\{\mathbf{e}_{uv}\}),$ (5)
where the gating coefficient $\gamma$ is dynamically learned by the network. This adaptive aggregation allows NeuralOM to efficiently handle complex multi-scale dynamics within a unified framework. For details on graph construction, encoder, and decoder, please refer to the Appendix.
Optimization and Inference
Objective.